OpenStudy (anonymous):

How would I solve for n in the problem: 1024=4(2)^n-1

OpenStudy (anonymous):

help me guys

OpenStudy (anonymous):

this is a geometric sequence

OpenStudy (anonymous):

you see, 1024 = 2^10

OpenStudy (anonymous):

simplified it equals 1024=8^n-1

OpenStudy (anonymous):

div both sides for 4, we have (2)^8 = (2)^n-1